OpenStudy (anonymous):

A die is rolled 12 times. Find the chance that the face with six spots appears once among the first 6 rolls, and once among the next 6 rolls.

OpenStudy (anonymous):

we should caluculate( 5/6)^10

OpenStudy (perl):

P( 6 spots is rolled exactly once on the first 6 throws & 6 spots is rolled exactly once on next 6 throws)

OpenStudy (perl):

probability of rolling a 6 is 1/6 let X = number of 6's rolled on first 6 rolls X is a binomial random variable n=6, p = 1/6 , P( X = 1 ) = 6C1 * (1/6)^1*(5/6)^5 Let Y = number of 6's rolled on next 6 rolls Y is a binomial random variable n=6, p = 1/6 , P( Y = 1 ) = 6C1 * (1/6)^1*(5/6)^5

OpenStudy (perl):

P(X = 6 & Y=6) = P(X=6)*P(Y=6) (since they are independent) = 6C1 * (1/6)^1*(5/6)^5*6C1 * (1/6)^1*(5/6)^5 = (5/6)^10

OpenStudy (kropot72):

@dinakar Yes, you should calculate (5/6)^10

OpenStudy (anonymous):

sry to ask u symbol(^) means multiplication

OpenStudy (perl):

it means x^2 = x*x 2^2 = 4

OpenStudy (perl):

(5/6)^10= (5/6)*(5/6)*(5/6)*(5/6)*(5/6)*(5/6)*(5/6)*(5/6)*(5/6)*(5/6)

OpenStudy (anonymous):

0.1615

OpenStudy (anonymous):

yes it was crct thanks